# Improving probabilistic error cancellation in the presence of non-stationary noise

### v1

@@ I Introduction
Contemporary quantum computers are not only noisy but they also exhibit non-stationarity. Non-stationary noise processes [20, 21, 22, 23, 24] in superconducting qubits [25] are well-studied. For example, state preparation and measurement (SPAM) fidelities have been observed to fluctuate significantly showing more than 25% deviation from their long-term average [26]. Similarly, the fidelity of CNOT gates have been noted to change by over 40% within similar time frames [27]. Moreover, the qubit relaxation times, known as $T_{1}$, have experienced fluctuations of up to 400% in just 30 minutes [28]. Likewise, de-phasing times, denoted as $T_{2}$, have been recorded to vary by over 50% within an hour [29, 30].
@@
Quantum error mitigation is a set of techniques that employs statistical tools from estimation theory to reduce the impact of noise in quantum computations without directly correcting the quantum state [18]. Such techniques can become vulnerable to errors stemming from over or under estimation of noise due to the presence of non-stationarity [35].
@@ III-A Uncorrelated parameters
$\mathcal{L}(f_{q})=f_{q}^{C_{0}[\mathcal{D}_{q}]}(1-f_{q})^{L-C_{0}[\mathcal{D}_{q}]}/B(\alpha_{q},\beta_{q})^{L}$ (6)
@@ III-B Correlated parameters
$\text{Pr}(p_{0},\cdots,p_{M-1})=\frac{\Gamma(\sum a_{i})}{\prod\limits_{0}^{M-1}\Gamma(a_{i})}\prod\limits_{0}^{M-1}p_{i}^{a_{i}-1}$ (8)
@@
with parameters $a_{i}^{\prime}=a_{i}+C_{v_{i}}(\mathcal{D})$. Upon marginalizing this joint density, the marginals follow a Beta distribution. For example, the random variable $p_{i}$ is Beta-distributed with parameters $\alpha_{i}=a_{i}^{\prime}$ and $\beta_{i}=-a_{i}^{\prime}+\sum\limits_{j=0}^{M-1}a_{j}^{\prime}$ [52].
The updated marginals from the Dirichlet distribution give us the evolving densities for each $p_{i}$, allowing us to calculate their time-varying means and variances. Specifically, the mean for $p_{i}$ updates to $a_{i}^{\prime}/\sum a_{i}^{\prime}$, and its variance updates to $a_{i}^{\prime}\left(\sum_{j}a_{j}^{\prime}-a_{i}^{\prime}\right)/\left(\sum_{j}a_{j}^{\prime}\right)^{2}/\left(1+\sum_{j}a_{j}^{\prime}\right)$.
@@ III-C Accuracy and stability
We next evaluate the performance of PEC in presence of non-stationary noise using the lens of accuracy and stability. A mean of a quantum observable $\mathcal{O}$ is represented as $\braket{\mathcal{O}}_{\textrm{x}}$ (where x labels the noise instance), while the observed mean after mitigation is denoted as $\braket{\mathcal{O}}_{\textrm{x}}^{\text{mit}}$, which equals $\braket{\mathcal{O}}$ only in the asymptotic limit of infinite samples and zero noise. When we want to say the mitigated noisy mean at a specific time $t$, we will use $\braket{\mathcal{O}}_{\textrm{x}}^{\text{mit}}(t)$.
@@
Note the dependence of $\eta_{{}_{\text{k}}}$ on the random variable x and time $t$.
When measuring $\mathcal{O}_{q}$ for qubit $q$, the qubit-wise accuracy metric is:
@@ IV Numerical Validation
A quantum circuit for the Bernstein-Vazirani algorithm is shown in Fig. 2. We conceptualize the circuit as having four principal layers ($\mathds{L}_{1},\mathds{L}_{2},\mathds{L}_{3}$, and $\mathds{L}_{4}$) separated by dashed vertical lines. Note that the total number of qubits needed by an $n-1$-bit secret string is $n$, comprising $n-1$ data qubits and one ancilla qubit. So, a 5-qubit implementation of the Bernstein-Vazirani algorithm has $n=5$ but the secret string length is $n-1=4$.
@@ IV-B Modeling circuit noise
In this expression, we use $\mathds{P}_{C}^{\prime},\mathds{P}_{T}^{\prime}\in\{\mathds{X},\mathds{Y},\mathds{Z}\}$ as the single-qubit Pauli operators excluding identity $\mathds{I}$, acting on the control (C) and target (T) qubits respectively. The sum is over all the single-qubit Pauli operators, excluding identity. The effect of the quantum noise channel $\mathcal{E}_{2}(\cdot)$ for layer 2, on a 5-qubit state, combines the identity channel $\mathcal{I}$ (which acts on qubits without CNOT connection) and $\mathcal{E}_{\text{CNOT}}$ (which acts on qubits linked by CNOT connections).
@@ IV-C1 CNOT noise
Using the requirement that the linear combination $\sum\eta_{{}_{k}}\tilde{\mathcal{G}}_{k}$ should equal the noiseless CNOT operation, we derive the PEC coefficients as follows.
Let $c_{0}=3/[\textrm{x}_{T}(1-\textrm{x}_{C})]^{2}+[\textrm{x}_{C}(1-\textrm{x}_{T})]^{2}-3c_{1}^{2}]$ and, $c_{1}=\textrm{x}_{C}+\textrm{x}_{T}-\textrm{x}_{C}\textrm{x}_{T}-1$.
When $\mathds{P}_{C}=\mathds{P}_{T}=\mathds{I}$,
$\eta_{{}_{\text{0}}}=c_{0}c_{1}.$
When $\mathds{P}_{C}=\mathds{I}$ and $\mathds{P}_{T}\in\{\mathds{X},\mathds{Y},\mathds{Z}\}$,
$\eta_{{}_{\text{k}}}=c_{0}\frac{\textrm{x}_{T}(1-\textrm{x}_{C})}{3},\;\;\;\;k\in\{1,2,3\}.$
When $\mathds{P}_{T}=\mathds{I}$ and $\mathds{P}_{C}\in\{\mathds{X},\mathds{Y},\mathds{Z}\}$,
$\eta_{{}_{\text{k}}}=c_{0}\frac{\textrm{x}_{C}(1-\textrm{x}_{T})}{3},\;\;\;\;k\in\{4,8,12\}.$
For all the remaining terms,
$\eta_{{}_{\text{k}}}=c_{0}\frac{\textrm{x}_{C}\textrm{x}_{T}}{9},\;\;\;\;k\in\{9,10,11,13,14,15\}.$
This yields the quasi-probability distribution as: $\{\mathds{Q}_{k}\}=\left\{|\eta_{{}_{0}}|/\gamma_{{}_{\text{CNOT}}},\cdots,|\eta_{{}_{15}}|/\gamma_{{}_{\text{CNOT}}}\right\}$ where $\gamma_{{}_{\text{CNOT}}}=|\eta_{{}_{0}}|+\cdots+|\eta_{{}_{15}}|.$
@@ IV-D Adaptive noise model
where $\Pi_{i}=\ket{i}\bra{i}$ | are the projection operators and $i\in\{00,01,10,11\}$. The probabilities are given by:
$\begin{split}\text{Pr}(00)=&f_{3}f_{4}(-1-\textrm{x}_{C}\textrm{x}_{T}+\textrm{x}_{C}+\textrm{x}_{T})\\ &+f_{3}(\textrm{x}_{C}\textrm{x}_{T}/2-\textrm{x}_{T}/2)\\ &+f_{4}(1-\textrm{x}_{T}-\textrm{x}_{C}/2+\textrm{x}_{C}\textrm{x}_{T}/2)\\ &-\textrm{x}_{C}\textrm{x}_{T}/4+\textrm{x}_{T}/2\\ \end{split}$ (29)
$\begin{split}\text{Pr}(01)=&f_{3}f_{4}(1-\textrm{x}_{C}-\textrm{x}_{T}+\textrm{x}_{C}\textrm{x}_{T})\\ &+f_{3}(-1-\textrm{x}_{C}\textrm{x}_{T}/2+\textrm{x}_{C}-\textrm{x}_{T}/2)\\ &+f_{4}(-1-\textrm{x}_{C}\textrm{x}_{T}/2+\textrm{x}_{C}+\textrm{x}_{T})\\ &+1+\textrm{x}_{C}\textrm{x}_{T}/4-\textrm{x}_{C}/2-\textrm{x}_{T}/2\\ \end{split}$ (30)
$\begin{split}\text{Pr}(10)=&f_{3}f_{4}(1-\textrm{x}_{T}-\textrm{x}_{C}+\textrm{x}_{C}\textrm{x}_{T})\\ &+f_{3}(\textrm{x}_{T}/2-\textrm{x}_{C}\textrm{x}_{T}/2)\\ &+f_{4}(\textrm{x}_{C}/2-\textrm{x}_{C}\textrm{x}_{T}/2)\\ &+\textrm{x}_{C}\textrm{x}_{T}/4\\ \end{split}$ (31)
$\begin{split}\text{Pr}(11)=&f_{3}f_{4}(-1+\textrm{x}_{C}+\textrm{x}_{T}-\textrm{x}_{C}\textrm{x}_{T})\\ &+f_{3}(1-\textrm{x}_{T}/2-\textrm{x}_{C}+\textrm{x}_{C}\textrm{x}_{T}/2)\\ &+f_{4}(\textrm{x}_{C}\textrm{x}_{T}/2-\textrm{x}_{C}/2)\\ &+\textrm{x}_{C}/2-\textrm{x}_{C}\textrm{x}_{T}/4\\ \end{split}$ (32)
In the last step, the updated PEC coefficients are obtained using Eqn. 27.
@@ IV-E Numerical simulation
For our 5-qubit circuit implementing the Bernstein-Vazirani algorithm, we used secret string $r=``1000"$. Thus, the qubit-wise mean of the $\mathds{Z}$ observable for the noiseless case is given by: $\braket{\mathcal{O}_{0}}=+1,\braket{\mathcal{O}_{1}}=+1,\braket{\mathcal{O}_{2}}=+1,\braket{\mathcal{O}_{3}}=-1,\braket{\mathcal{O}_{4}}=+1$, qubit 4 being the ancilla.
@@ V Experimental Testing
Figure 8: The figure comprises two plots, each depicting four graphs: (1) "No mitigation" presents raw Bernstein-Vazirani metrics without error mitigation, (2) "ROEM" shows metrics after readout error mitigation, with constant SPAM noise parameters, (3) "Non-adaptive PEC" displays metrics for PEC, and (4) "Adaptive PEC" exhibits metrics for PEC with adaptive noise mitigation. The x-axis denotes intra-calibration time-stamps (UTC) for Jan 15. Plot (a) illustrates the time-varying accuracy metric from Eqn. 16. It demonstrates the adaptive PEC’s 42% accuracy improvement over non-adaptive PEC. Plot (b) shows the time-varying stability metric from Eqn. 19. It illustrates the adaptive PEC’s 60% stability enhancement compared to non-adaptive PEC. These plots underscore the significant impact of non-stationary noise on PEC resilience. Due to PEC’s lengthy completion time (a couple of hours), adaptive methods are able to handle non-stationary noise conditions better.
@@ V-A Non-stationary noise estimates
Fig. 6 illustrates how noise in the quantum computer changed over time. The blue line in plot (a) shows the depolarizing parameter for the target qubit of the CNOT gate, while the black line represents the depolarizing parameter for the control qubit. Plot (b) shows five lines, each representing the SPAM fidelity for the register elements. The x-axis denotes intra-calibration timestamps.
The graph shows periods where the noise levels in the depolarizing parameter for qubit 3 (the control qubit) are steady, notably between 2:00 am and 7:30 am, contrasting with times of significant fluctuation, as observed between 9:00 am and 2:30 pm. The depolarizing parameter fluctuates between 1% and 4% for qubit 4 (the target qubit) and between 1% and 3% for qubit 3, both peaking sharply at 10:53 am.
Qubit 4 experiences the most significant impact from SPAM noise, with its values fluctuating between 0.99 and 0.94, a notable range given the sensitivity of PEC to accurate noise estimations. In contrast, Qubit 3 maintains a consistent SPAM fidelity throughout the same period, indicated by small error margins and a steady average value of 0.99. Meanwhile, Qubit 2 demonstrates a gradual drift in its values, starting from below 0.94 and rising to 0.96. This progressive change suggests a systematic, non-random trend that might be rectifiable with bias shift corrections. However, such patterns are not uniform across the entire register, implying the necessity to consider non-stationary statistics for modeling the system. Conducting experiments in times of significant non-stationary activity, like from 7:30 am to 2:30 pm, lead to more unstable outcomes when using non-adaptive PEC.

### v2

@@ I Introduction
Contemporary quantum computers are not only noisy but they also exhibit non-stationarity. Non-stationary noise processes [20, 21, 22, 23, 24] in superconducting qubits [25] are well-studied. For example, state preparation and measurement (SPAM) fidelities have been observed to fluctuate significantly, showing more than 25% deviation from their long-term average [26]. Similarly, the fidelity of CNOT gates have been noted to change by over 40% within similar time frames [27]. Moreover, the qubit relaxation time ($T_{1}$) has been observed to experience fluctuations of up to 400% in just 30 minutes [28]. Likewise, de-phasing time ($T_{2}$) has been recorded to vary by over 50% within an hour [29, 30].
@@
Quantum error mitigation is a set of techniques that employs statistical tools from estimation theory to reduce the impact of noise in quantum computations without directly correcting the quantum state [18]. Such techniques can become vulnerable to errors stemming from over or under estimation of noise due to the presence of non-stationarity [35, 36].
@@ III-A Uncorrelated parameters
$\mathcal{L}(f_{q})=f_{q}^{C_{0}[\mathcal{D}_{q}]}(1-f_{q})^{L-C_{0}[\mathcal{D}_{q}]}/B^{L}(\alpha_{q},\beta_{q})$ (6)
@@ III-B Correlated parameters
$\text{Pr}(p_{0},\cdots,p_{M-1})=\frac{\Gamma(\sum a_{i})}{\prod\limits_{i=0}^{M-1}\Gamma(a_{i})}\prod\limits_{i=0}^{M-1}p_{i}^{a_{i}-1}$ (8)
@@
with parameters
$a_{i}^{\prime}=a_{i}+C_{v_{i}}(\mathcal{D}).$ (11)
Upon marginalizing this joint density, the marginals follow a Beta distribution. For example, the random variable $p_{i}$ is Beta-distributed with parameters [53]:
$\begin{split}\alpha_{i}=&a_{i}^{\prime}\\ \beta_{i}=&-a_{i}^{\prime}+\sum\limits_{j=0}^{M-1}a_{j}^{\prime}\end{split}$ (12)
The updated marginals from the Dirichlet distribution give us the evolving densities for each $p_{i}$, allowing us to calculate their time-varying means and variances. Specifically, the mean for $p_{i}$ updates to:
$\mathds{E}(p_{i})=\frac{a_{i}^{\prime}}{\sum\limits_{j}a_{j}^{\prime}}$ (13)
and its variance updates to:
$\text{Var}(p_{i})=\frac{a_{i}^{\prime}\left(\sum\limits_{j}a_{j}^{\prime}-a_{i}^{\prime}\right)}{\left(\sum\limits_{j}a_{j}^{\prime}\right)^{2}\left(1+\sum\limits_{j}a_{j}^{\prime}\right)}.$ (14)
@@ III-C Accuracy and stability
We next evaluate the performance of PEC in presence of non-stationary noise using the lens of accuracy and stability. A mean of a quantum observable $\mathcal{O}$ is represented as $\braket{\mathcal{O}}_{\textrm{x}}$ (where x labels the noise instance), while the observed mean after mitigation is denoted as $\braket{\mathcal{O}}_{\textrm{x}}^{\text{mit}}$, which equals $\braket{\mathcal{O}}$ only in the asymptotic limit of infinite samples and zero noise. To denote the mitigated noisy mean at a specific time $t$, we use $\braket{\mathcal{O}}_{\textrm{x}}^{\text{mit}}(t)$.
@@
Note the dependence of $\eta_{{}_{\text{k}}}$ on the random variable x and time $t$. When measuring $\mathcal{O}_{q}$ for qubit $q$, the qubit-wise accuracy metric is:
@@ IV Numerical Validation
A quantum circuit for the Bernstein-Vazirani algorithm is shown in Fig. 2. We conceptualize the circuit as having four principal layers ($\mathds{L}_{1},\mathds{L}_{2},\mathds{L}_{3}$, and $\mathds{L}_{4}$) separated by dashed vertical lines. Note that the total number of qubits needed by a $(n-1)$-bit secret string is $n$, comprising $n-1$ data qubits and one ancilla qubit. So, a 5-qubit implementation of the Bernstein-Vazirani algorithm has $n=5$ but the secret string length is $n-1=4$.
@@ IV-B Modeling circuit noise
In this expression, we use $\mathds{P}_{C}^{\prime},\mathds{P}_{T}^{\prime}\in\{\mathds{X},\mathds{Y},\mathds{Z}\}$ as the single-qubit Pauli operators excluding identity $\mathds{I}$, acting on the control (C) and target (T) qubits respectively. The sum is over all the single-qubit Pauli operators, excluding identity. Note that the effect of the quantum noise channel $\mathcal{E}_{2}(\cdot)$ for layer 2, on a 5-qubit state, combines the identity channel $\mathcal{I}$ (which acts on qubits without CNOT connection) and $\mathcal{E}_{\text{CNOT}}$ (which acts on qubits linked by CNOT connections).
@@ IV-C1 CNOT noise
Using the requirement that the linear combination $\sum\eta_{{}_{k}}\tilde{\mathcal{G}}_{k}$ should equal the noiseless CNOT operation, we derive the PEC coefficients as follows. When $\mathds{P}_{C}=\mathds{P}_{T}=\mathds{I}$, $\eta_{{}_{\text{0}}}=c_{0}c_{1}$. When $\mathds{P}_{C}=\mathds{I}$ and $\mathds{P}_{T}\in\{\mathds{X},\mathds{Y},\mathds{Z}\}$, $\eta_{{}_{\text{k}}}=c_{0}\textrm{x}_{T}(1-\textrm{x}_{C})/3,\;\;\;\;k\in\{1,2,3\}$. When $\mathds{P}_{T}=\mathds{I}$ and $\mathds{P}_{C}\in\{\mathds{X},\mathds{Y},\mathds{Z}\}$, $\eta_{{}_{\text{k}}}=c_{0}\textrm{x}_{C}(1-\textrm{x}_{T})/3,\;\;\;\;k\in\{4,8,12\}$. For all the remaining terms, $\eta_{{}_{\text{k}}}=c_{0}\textrm{x}_{C}\textrm{x}_{T}/9,\;\;\;\;k\in\{9,10,11,13,14,15\}$. Here, $c_{0}=3/[\textrm{x}_{T}(1-\textrm{x}_{C})]^{2}+[\textrm{x}_{C}(1-\textrm{x}_{T})]^{2}-3c_{1}^{2}]$ and, $c_{1}=\textrm{x}_{C}+\textrm{x}_{T}-\textrm{x}_{C}\textrm{x}_{T}-1$. This yields the quasi-probability distribution as:
$\{\mathds{Q}_{k}\}=\left\{|\eta_{{}_{0}}|/\gamma_{{}_{\text{CNOT}}},\cdots,|\eta_{{}_{15}}|/\gamma_{{}_{\text{CNOT}}}\right\}$ (28)
where $\gamma_{{}_{\text{CNOT}}}=|\eta_{{}_{0}}|+\cdots+|\eta_{{}_{15}}|.$
@@ IV-D Adaptive noise model
where $\Pi_{i}=\ket{i}\bra{i}$ | are the projection operators and $i\in\{00,01,10,11\}$. For our running example, the probabilities are given by:
$\begin{split}\text{Pr}(00)=&f_{3}f_{4}(-1-\textrm{x}_{C}\textrm{x}_{T}+\textrm{x}_{C}+\textrm{x}_{T})\\ &+f_{3}(\textrm{x}_{C}\textrm{x}_{T}/2-\textrm{x}_{T}/2)\\ &+f_{4}(1-\textrm{x}_{T}-\textrm{x}_{C}/2+\textrm{x}_{C}\textrm{x}_{T}/2)\\ &-\textrm{x}_{C}\textrm{x}_{T}/4+\textrm{x}_{T}/2\\ \text{Pr}(01)=&f_{3}f_{4}(1-\textrm{x}_{C}-\textrm{x}_{T}+\textrm{x}_{C}\textrm{x}_{T})\\ &+f_{3}(-1-\textrm{x}_{C}\textrm{x}_{T}/2+\textrm{x}_{C}-\textrm{x}_{T}/2)\\ &+f_{4}(-1-\textrm{x}_{C}\textrm{x}_{T}/2+\textrm{x}_{C}+\textrm{x}_{T})\\ &+1+\textrm{x}_{C}\textrm{x}_{T}/4-\textrm{x}_{C}/2-\textrm{x}_{T}/2\\ \text{Pr}(10)=&f_{3}f_{4}(1-\textrm{x}_{T}-\textrm{x}_{C}+\textrm{x}_{C}\textrm{x}_{T})\\ &+f_{3}(\textrm{x}_{T}/2-\textrm{x}_{C}\textrm{x}_{T}/2)\\ &+f_{4}(\textrm{x}_{C}/2-\textrm{x}_{C}\textrm{x}_{T}/2)\\ &+\textrm{x}_{C}\textrm{x}_{T}/4\\ \text{Pr}(11)=&f_{3}f_{4}(-1+\textrm{x}_{C}+\textrm{x}_{T}-\textrm{x}_{C}\textrm{x}_{T})\\ &+f_{3}(1-\textrm{x}_{T}/2-\textrm{x}_{C}+\textrm{x}_{C}\textrm{x}_{T}/2)\\ &+f_{4}(\textrm{x}_{C}\textrm{x}_{T}/2-\textrm{x}_{C}/2)\\ &+\textrm{x}_{C}/2-\textrm{x}_{C}\textrm{x}_{T}/4\\ \end{split}$ (34)
In the last step, the updated PEC coefficients are obtained using Eqn. 32.
@@ IV-E Numerical simulation
For our 5-qubit circuit implementing the Bernstein-Vazirani algorithm, we used secret string $r=1000$. Thus, the qubit-wise mean of the $\mathds{Z}$ observable for the noiseless case is given by: $\braket{\mathcal{O}_{0}}=+1,\braket{\mathcal{O}_{1}}=+1,\braket{\mathcal{O}_{2}}=+1,\braket{\mathcal{O}_{3}}=-1,\braket{\mathcal{O}_{4}}=+1$, qubit 4 being the ancilla.
@@ V Experimental Testing
Figure 8: The figure comprises two plots, each depicting four graphs: (1) "No mitigation" presents raw Bernstein-Vazirani metrics without error mitigation, (2) "ROEM" shows metrics after readout error mitigation, with constant SPAM noise parameters, (3) "Non-adaptive PEC" displays metrics for PEC, and (4) "Adaptive PEC" exhibits metrics for PEC with adaptive noise mitigation. The x-axis denotes intra-calibration time-stamps (UTC) for Jan 15. Plot (a) illustrates the time-varying accuracy metric from Eqn. 20. It demonstrates the adaptive PEC’s 42% accuracy improvement over non-adaptive PEC. Plot (b) shows the time-varying stability metric from Eqn. 23. It illustrates the adaptive PEC’s 60% stability enhancement compared to non-adaptive PEC. These plots underscore the significant impact of non-stationary noise on PEC resilience. Due to PEC’s lengthy completion time (a couple of hours), adaptive methods are able to handle non-stationary noise conditions better.
@@ V-A Non-stationary noise estimates
Fig. 6 illustrates how noise in the quantum computer changed over time. The blue line in plot (a) shows the depolarizing parameter for the target qubit of the CNOT gate, while the black line represents the depolarizing parameter for the control qubit. Plot (b) shows five lines, each representing the SPAM fidelity for the register elements. The x-axis denotes intra-calibration timestamps. The graph shows periods where the noise levels in the depolarizing parameter for qubit 3 (the control qubit) are steady, notably between 2:00 am and 7:30 am, contrasting with times of significant fluctuation, as observed between 9:00 am and 2:30 pm. The depolarizing parameter fluctuates between 1% and 4% for qubit 4 (the target qubit) and between 1% and 3% for qubit 3, both peaking sharply at 10:53 am.
Since the probabilities of observing the computational basis states on the left-hand side in Eqn. 33 are functions of time, we expect the SPAM fidelities to also vary over time. However, from Fig. 6 (b), the value for qubit 3 appears almost constant. This qubit exhibits less non-stationarity compared to other qubits, suggesting greater reliability with respect to SPAM noise. However, these estimates are influenced by grid discretization used in the mean square error minimization procedure for solving the non-linear equations to find the best-fit. Our grid for SPAM fidelity had a spacing of $0.01$ in the range $[0,1]$. Our optimizer consistently produced an estimate for $f_{3}$ as $0.99$ across all the time-stamps, which is near the upper limit of the permissible range. We have not analyzed the impact of discretization granularity on degree of non-stationarity of the solution space.
Qubit 4 experiences the most significant impact from SPAM noise, with its values fluctuating between 0.99 and 0.94, a notable range given the sensitivity of PEC to accurate noise estimations. Meanwhile, Qubit 2 demonstrates a gradual drift in its values, starting from below 0.94 and rising to 0.96. This progressive change suggests a systematic, non-random trend that might be rectifiable with bias shift corrections. However, such patterns are not uniform across the entire register, implying the necessity to consider non-stationary statistics for modeling the system. Conducting experiments in times of significant non-stationary activity, like from 7:30 am to 2:30 pm, lead to more unstable outcomes when using non-adaptive PEC.